OpenStudy (anonymous):

the probability that it would snow on sunday is 3/5 what is the probabilty that it would snow on monday, if it snowed on sunday ?

OpenStudy (anonymous):

there must be some more information

OpenStudy (anonymous):

ooh sorry the probability that it would snow on sunday is 3/5 . the probability that it will snow on both sunday and monday is 3/10 . what is the probabilty that it would snow on monday, if it snowed on sunday

OpenStudy (anonymous):

\[P(A|B)=\frac{P(A\cap B)}{P(B)}\]

OpenStudy (anonymous):

\[P(A|B)=\frac{\frac{3}{10}}{\frac{3}{5}}=\frac{3}{10}\times \frac{5}{3}=\frac{1}{2}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome